# Psycholinguistic Analyses in Software Engineering Text: A Systematic Literature Review

### v1

@@ 5.1 RQ1: How was LIWC employed in SE studies, and for what specific activities and purposes?
Team Management Group Communication Analysis in Teams D Pronouns, Cognitive, Work and Achievement, Leisure, Social and Positive, Negative Language S26, S32, S34, S36
Team Management Group Communication Analysis in Teams D Cognitive, Work and Achievement, Leisure, Social and Positive, Negative Language S29, S33
Team Management Group Communication Analysis in Teams D Achieve, Affect, Discrep, Funct, Hear, Humans, Leisure, Negative Emotion, Numerals, Past, Present, QuestionMark, Social S18
Team Management Group Communication Analysis in Teams D Work and achievement, negative, cognitive, social and positive S17
Team Management Group Communication Analysis in Teams D Pronouns, Work and Achievement, Cognitive, Leisure, Social and Emotions S10
Team Management Emotion Detection D Positive affect (Positive Emotion), Negative affect (Negative Emotion, Sadness, Anger, Anxiety, Inhibition) S30
Quality assurance process User Input Analysis D Negative Emotion S38
Quality assurance process LLM Evaluation D Affective Process, Cognitive Process, Perceptual Process, Informal language S43
Team Management Emotion Detection D Categories Not Specified S37
Team Management Group Communication Analysis in Teams D Linguistic process, Linguistic categories, Psychological Processes, relativity, current concerns, spoken categories, and punctuation S6
Team Management Group Communication Analysis in Teams D Pronouns, Cognitive, Work and Achievement, Emotions S5
Team Management Emotion Detection D Anger S28
Team Management Group Communication Analysis in Teams D prop(focusfuture), prop(we), prop(posemo), prop(negemo), prop(certain), prop(tentative) S9
@@
Team Management Personality Detection I Categories correlated with personality traits in [49] S2, S20, S23
Quality Assurance Process User Input Analysis I Negative Emotion S15, S22, S40
Quality Assurance Process Insider Threat Detection I Categories Not Specified S41
Stakeholder Needs and Requirements Definition Identifying Developer Preferences I Visual, Auditive, “Kinaesthetic” S42
Team Management Emotion Detection I Categories Not Specified S12, S13
Team Management Personality Detection I Categories correlated with personality traits in [50] and [51] S19
Team Management Emotion Detection I Anger S14
Team Management Sentiment Analysis I Categories Not Specified S3
Team Management Personality Detection I Categories correlated with personality traits in [50] S7
Team Management Group Communication Analysis in Teams I Categories Not Specified S4
Knowledge Management Predicting Deletion of Stack Overflow Post I Personal Pronouns, Pronouns, Space words, Relativity words, Inclusive words, Cognitive Process words, Social words, 1st person singular pronouns, Function Words, Conjunctions, Prepositions S1
Knowledge Management Predicting Deletion of Stack Overflow Post I Personal Pronouns, Pronouns, Space, Relativity, Inclusive, Cognitive Process, Social, 1st Person Singular Pronouns, Function Words, Conjunctions, Prepositions S39
Team Management Group Communication Analysis in Teams I Personal pronouns, Impersonal Pronouns, Articles, Prepositions, Auxiliary verbs, Common adverbs, Conjunctions, Negations, Analytical Thinking, Clout, Authentic, and Emotional Tone. S24
Team Management Personality Detection I Categories correlated with personality traits in [50], Categories used in Personality Recognizer tool  [^1] http://s3.amazonaws.com/mairesse/research/personality/recognizer.html S8
Team Management Group Communication Analysis in Teams D & I Standard WC Pronouns, Cognitive, Relative, Optimism, References to Time, Future tense verbs, Social Processes, Categories correlated with personality traits in [49] S27
Team Management Personality Detection D & I Categories correlated with personality traits in [49], Cognitive, Work and Achievement, Leisure, Social and Positive, Negative Language S35
Team Management Argument Extraction D & I Categories Not Specified S11
Among all the included studies in our survey, 19 used LIWC to directly analyze text. As the name suggests, direct use of LIWC refers to the cases in which researchers used LIWC’s outputs directly to interpret and gain insights from the textual data without any further processing. These use-cases frequently encompass tasks such as detecting developer emotions (e.g., anger, negative emotion, and anxiety) [S13, S18, S28, S30, S37] using LIWC categories like affect, emotions, etc. Other tasks included assessing characteristics, such as goal orientation and sociability [S5, S10, S23, S26] and evaluating the linguistic features of LLM responses to Stack Overflow questions.
On the other hand, 21 studies employed LIWC’s output indirectly, using it to develop other algorithms or models or to extract insights that necessitated further processing of LIWC’s outputs. The most common method of using LIWC indirectly was to uncover developers’ personalities [S2, S8, S9, S19, S27, S35]. Indeed, prior research has demonstrated multiple methods of inferring text authors’ personalities using LIWC. Pennebaker & King [49], Yarkoni et al. [50], and Goldbeck et al. [51] all have proposed methods of leveraging LIWC for measuring the personality traits of an author based on the widely used Big-Five Factor personality model [52]. However, it should be noted that while the original studies that employ LIWC categories to measure personality traits have received significant attention [49, 50, 51], their focus has predominantly centered on general rather than domain-specific text, such as software engineering-related content. Next, we describe some of the other distinctive indirect applications of LIWC in software engineering-related research.
@@
LIWC has also been in the development of other psychometric tools. In another work, Santos et al. present NEUROMINER, a psychometric tool based on LIWC, that identifies developers’ preferred representational systems [S25]. Representational systems refer to the ways in which the human brain stores and processes information. Often within the NLP domain, these systems are recognized as Visual, Auditive, or Kinaesthetic [54]. Knowing the developers’ preferred representational systems can help us better understand the most effective methods of communication and knowledge sharing. NEUROMINER extracts LIWC values from an input text, and performs ANOVA (analysis of variance) to detect the preferred representational systems of the top four Apache committers. The authors use Visual, Auditive, and Kinaesthetic categories of LIWC to decide which of these representational models a developer finds more comfortable.
@@
Additionally, [S19] and [S8] examine LIWC’s ability to infer developers’ personalities and compare the tool with the now-discontinued LIWC-based tool, IBM Personality Insights  [^2] https://watson-developer-cloud.github.io/node-sdk/master/classes/personalityinsightsv3.html. While [S19] finds the two tools to be consistent in what results they achieve, [S8] points out a noticeable level of disagreement between LIWC and IBM Personality Insights.
@@
[backgroundcolor=lightgray!20,topline=false,leftline=false,rightline=false,bottomline=false] RQ1 Summary: LIWC has been applied to a variety of SE tasks to understand the nuances of developer learning, collaboration and communication. This includes emotion and sentiment analysis, personality detection, quality assessment of forum posts, investigation of cultural differences in distributed teams, among others.
@@ 5.2 RQ2: What types of textual data were analyzed using LIWC in SE studies?
Interactions on project management platforms IBM rational jazz S2, S17, S26, S29, S31, S32, S33, S34, S36
@@
Most researchers focused on analyzing the communications of developers on project management platforms, i.e., IBM Rational Jazz, GitHub, and Jira. In 9 of the 43 studies, IBM Rational Jazz was used as the primary source of data. In addition to hosting developers’ communications, IBM Rational Jazz provides environments for both development and tracking the progress of the project. These 9 papers use LIWC to study developers’ comments in Jazz repositories over time, and to explore various aspects of developer interactions and social dynamics, including the impact of leadership and knowledge sharing on project success [S26, S31].
@@ 5.3 RQ3: What Behavioral Software Engineering (BSE) concepts were studied using LIWC?
Understanding how LIWC has been used to study BSE concepts reveals the tools’ effectiveness in capturing critical human factors that directly influence software engineering practices and outcomes. Answering this research question contributes to assessing LIWC’s role in advancing behavioral research within SE. Next we discuss in detail the studied behavioral software engineering concepts using LIWC in our 43 selected studies.
To find out whether or not LIWC has been used to conduct software engineering-related studies tackling different BSE concepts, we have identified all the BSE concepts studied in each of our selected studies. Table LABEL:tab:bse illustrates the definitions of the BSE concepts that have been borrowed from  [26] and identified in one or more of our selected 40 studies.
@@
Organizational Climate “Is defined as the recurring patterns of behavior, attitudes and feelings that characterize life in the organization (Isaksen et al., 2007; Denison, 1996).” [26, 68, 69] S4, S5, S7, S12, S14, S24, S27, S28, S29, S30, S31, S32, S33, S34, S36
Positive psychology “Positive psychology is the branch of psychology that study the strengths and virtues that enable individuals and communities to thrive (Seligman and Csikszentmihalyi, 2000).” [26, 70] [S3, S17, S18, S23, S26, S27, S28, S29, S30, S31, S33, S35
Personality “According to Feist and Feist (Feist, 1994) personality is a pattern of relatively permanent traits and unique characteristics that give both consistency and individuality to a person’s behavior.” [26, 71] S2, S7, S8, S17, S19, S20, S27, S35
@@
Motivation “According to Pardee (Pardee, 1990) motivation can be defined as those forces within an individual that push or propel him to satisfy basic needs. It is what prompts a person to act in a certain way or at least develop an inclination for specific behavior.” [26, 74] S23, S27
@@
Group dynamics “Is a system of behaviors and psychological processes that occurs within a group or between groups (Forsyth, 2009).” [26, 78] S24
Group social identity “Defined by Hogg and Vaughan as an self-concept derived from perceived membership of social groups, i.e. the aspects of a person that are defined in terms of his or her group memberships (Hogg and Vaughan, 2002).” [26, 79] S10
Social value orientation “The stable preferences for certain patterns of outcomes for oneself and others, e.g. how to allocate resources (e.g. money) between your self and another person (Griesinger and Livingston, 1973).” [26, 80] S10
@@
Organizational Climate, Positive Psychology, and Personality represent the most extensively explored concepts within the collected body of literature. Among our subjects of study, the least commonly studied BSE concepts were Social Value Orientation, Stress, and Group Dynamics.
In line with Lenberg et al.’s results, indicating Communication and Personality to be the most studied BSE concepts, we found both of these concepts to be frequently examined in papers employing LIWC in software engineering. However, contrary to the observations by Lenberg et al., we found that Organizational Climate and Positive Psychology also to be be frequently examined and studied in a number of papers.
@@
[backgroundcolor=lightgray!20,topline=false,leftline=false,rightline=false,bottomline=false] RQ3 Summary: We present all the BSE concepts studied with the use of LIWC in Table LABEL:tab:bse. The most commonly examined concepts include Communication, Organizational Climate, Positive Psychology, and Personality, while concepts such as Social Value Orientation, Stress, and Group Dynamics, are less commonly studied.
@@ 5.4 RQ4: How often and in what way has LIWC been evaluated in a software engineering context?
A notable observation among the papers that used LIWC is that the majority, specifically 26 of 43 papers, did not formally evaluate LIWC’s output (see Table 8). These papers relied on LIWC for various purposes and accepted the resulting psychometric measures without any evaluation. In these studies, LIWC’s output was either used for qualitatively analyzing the text or was combined with other data to facilitate subsequent analysis. Some of these studies have used the combination of LIWC’s output and other data points to train machine learning models for tasks such as the prediction of the low quality Stack Overflow posts [S1, S39] or the identification of different arguments within a discussion i.e, argument extraction [S11]. These papers mainly justified their decision based on the established reputation of LIWC as a widely recognized and reliable psychometric text analysis tool, cited by thousands of studies  [22].
@@
One study, [S9], performed a comparison of LIWC’s sentiment analysis performance with another model, NLTK VADER, and found them to be on par. However, this was not a formal evaluation of LIWC itself, as it simply compared the performance between the two tools without directly assessing LIWC’s internal accuracy or outputs in isolation. Despite this, we still wanted to distinguish it as a valuable comparison for understanding how LIWC performs relative to other sentiment analysis tools.

### v2

@@ 5.1.2 Direct and Indirect Use of LIWC
Team Management Group Communication Analysis in Teams D Pronouns, Cognitive, Work and Achievement, Leisure, Social and Positive, Negative Language S26, S31, S32, S34, S36
Group Communication Analysis in Teams D Cognitive, Work and Achievement, Leisure, Social and Positive, Negative Language S29, S33
Identifying Developer Preferences I Visual, Auditive, “Kinaesthetic” S25
Group Communication Analysis in Teams D Achieve, Affect, Discrep, Funct, Hear, Humans, Leisure, Negative Emotion, Numerals, Past, Present, QuestionMark, Social S18
Group Communication Analysis in Teams D Work and achievement, negative, cognitive, social and positive S17
Group Communication Analysis in Teams D Linguistic process, Linguistic categories, Psychological Processes, relativity, current concerns, spoken categories, and punctuation S6
Group Communication Analysis in Teams D Pronouns, Cognitive, Work and Achievement, Emotions S5
Group Communication Analysis in Teams D Pronouns, Work and Achievement, Cognitive, Leisure, Social and Emotions S10
Group Communication Analysis in Teams D Achievement, Work, Positive, Negative, Insight, Social, prop(I) S37
Group Communication Analysis in Teams D prop(focusfuture), prop(we), prop(posemo), prop(negemo), prop(certain), prop(tentative) S9
Group Communication Analysis in Teams I Categories Not Specified S4
Group Communication Analysis in Teams I Personal pronouns, Impersonal Pronouns, Articles, Prepositions, Auxiliary verbs, Common adverbs, Conjunctions, Negations, Analytical Thinking, Clout, Authentic, and Emotional Tone. S24
Group Communication Analysis in Teams D & I Standard WC Pronouns, Cognitive, Relative, Optimism, References to Time, Future tense verbs, Social Processes, Categories correlated with personality traits in [60] S27
Personality Detection I Categories correlated with personality traits in [60] S2, S20, S23
Personality Detection I Categories correlated with personality traits in [30] and [61] S19
Personality Detection I Categories correlated with personality traits in [30] S7
Personality Detection I Categories correlated with personality traits in [30], Categories used in Personality Recognizer tool  [^3] http://s3.amazonaws.com/mairesse/research/personality/recognizer.html S8
Personality Detection D & I Categories correlated with personality traits in [60], Cognitive, Work and Achievement, Leisure, Social and Positive, Negative Language S35
Emotion Detection D Positive affect (Positive Emotion), Negative affect (Negative Emotion, Sadness, Anger, Anxiety, Inhibition) S30
Emotion Detection D Anger S28
Emotion Detection I Categories Not Specified S12, S13
Emotion Detection I Anger S14
Sentiment Analysis I Categories Not Specified S3
Argument Extraction D & I Categories Not Specified S11
Quality Assurance Predicting Deletion of Stack Overflow Post I Personal Pronouns, Pronouns, Space, Relativity, Inclusive, Cognitive Process, Social, 1st Person Singular Pronouns, Function Words, Conjunctions, Prepositions S1, S39
LLM Evaluation D Affective Process, Cognitive Process, Perceptual Process, Informal language S43
Insider Threat Detection I Categories Not Specified S41
@@
5.1.3 Notable Applications of LIWC
Through answering RQ1, we also aim to better understand the types of LIWC applications within the SE research. We pinned down the purposes LIWC has fulfilled in each study. Here, in more detail, we discuss some notable applications of LIWC in SE research.
@@ 5.1.3 Notable Applications of LIWC
LIWC has also been used in the development of other psychometric tools. In another work, Santos et al. present NEUROMINER, a psychometric tool based on LIWC, that identifies developers’ preferred representational systems [S25]. Representational systems refer to the ways in which the human brain stores and processes information. Often within the NLP domain, these systems are recognized as Visual, Auditive, or Kinaesthetic [64]. Knowing the developers’ preferred representational systems can help us better understand the most effective methods of communication and knowledge sharing. NEUROMINER extracts LIWC values from an input text, and performs ANOVA (analysis of variance) to detect the preferred representational systems of the top four Apache committers. The authors use Visual, Auditive, and Kinaesthetic categories of LIWC to decide which of these representational models a developer tends to prefer.
@@
Additionally, [S19] and [S8] examine LIWC’s ability to infer developers’ personalities and compare the tool with the now-discontinued LIWC-based tool, IBM Personality Insights  [^4] https://watson-developer-cloud.github.io/node-sdk/master/classes/personalityinsightsv3.html. While [S19] finds the two tools to produce consistent results, [S8] points out a noticeable level of disagreement between LIWC and IBM Personality Insights.
@@
[backgroundcolor=lightgray!20,topline=false,leftline=false,rightline=false,bottomline=false] RQ1 Summary: LIWC has been directly and indirectly utilized in a variety of SE tasks to understand the nuances of developer learning, collaboration, and communication. The most common uses include personality detection and emotion analysis. Other applications include quality assessment of forum posts, investigation of cultural differences in distributed teams, and argument extraction.
@@ 5.2 RQ2: What types of textual data were analyzed using LIWC in SE studies?
Interactions on project management platforms IBM rational jazz S2, S17, S26, S29, S31, S32, S33, S34, S36, S37
@@
Most researchers focused on analyzing the communications of developers on project management platforms, i.e., IBM Rational Jazz, GitHub, and Jira. In 10 of the 43 studies, IBM Rational Jazz was used as the primary source of data. In addition to hosting developers’ communications, IBM Rational Jazz provides environments for both development and tracking the progress of the project. These 10 papers use LIWC to study developers’ comments in Jazz repositories over time, and to explore various aspects of developer interactions and social dynamics, including the impact of leadership and knowledge sharing on project success [S26, S31]. Figure 2 provides a visual overview of the distribution of data sources across the studies we reviewed.
Figure 2: Distribution of Data Sources Used in SE Studies Employing LIWC. The figure highlights the prevalence of project management platforms and Q&A forums as dominant sources of analyzed text.
@@ 5.3 RQ3: What Behavioral Software Engineering (BSE) concepts were studied using LIWC?
Understanding how LIWC has been used to study BSE concepts reveals the tool’s effectiveness in capturing critical human factors that directly influence software engineering practices and outcomes. Answering this research question contributes to assessing LIWC’s role in advancing behavioral research within SE. Next we discuss in detail the studied behavioral software engineering concepts using LIWC in our 43 selected studies.
To find out whether or not LIWC has been used to conduct software engineering-related studies tackling different BSE concepts, we have identified all the BSE concepts studied in each of our selected studies. Table LABEL:tab:bse illustrates the definitions of the BSE concepts that have been borrowed from  [33] and identified in one or more of our selected 43 studies.
@@
Organizational Climate “Is defined as the recurring patterns of behavior, attitudes and feelings that characterize life in the organization (Isaksen et al., 2007; Denison, 1996).” [33, 76, 77] S4, S5, S7, S12, S14, S24, S27, S28, S29, S30, S31, S32, S33, S34, S36, S37
Motivation “According to Pardee (Pardee, 1990) motivation can be defined as those forces within an individual that push or propel him to satisfy basic needs. It is what prompts a person to act in a certain way or at least develop an inclination for specific behavior.” [33, 32] S7, S9, S12, S17, S21, S22, S26, S31, S33, S36, S37, S38, S40
Positive psychology “Positive psychology is the branch of psychology that study the strengths and virtues that enable individuals and communities to thrive (Seligman and Csikszentmihalyi, 2000).” [33, 78] [S3, S17, S18, S23, S26, S27, S28, S29, S30, S31, S33, S35, S37
Personality “According to Feist and Feist (Feist, 1994) personality is a pattern of relatively permanent traits and unique characteristics that give both consistency and individuality to a person’s behavior.” [33, 79] S2, S7, S8, S17, S19, S20, S27, S35, S37
@@
Group dynamics “Is a system of behaviors and psychological processes that occurs within a group or between groups (Forsyth, 2009).” [33, 81] S2, S4, S6, S10, S11, S24, S28
@@
Group social identity “Defined by Hogg and Vaughan as an self-concept derived from perceived membership of social groups, i.e., the aspects of a person that are defined in terms of his or her group memberships (Hogg and Vaughan, 2002).” [33, 86] S10
Risk taking “Is the tendency to engage in potential harmful behavior that, at the same time, could have some kind of positive outcome (March and Shapira, 1987).” [33, 87] S41
@@
Organizational Climate, Motivation, Positive Psychology, and Personality represent the most extensively explored concepts within the collected body of literature. Among our subjects of study, the least commonly studied BSE concepts were Risk taking, Stress, and Group Dynamics.
In line with Lenberg et al.’s results, indicating Communication and Personality to be among the most studied BSE concepts, we found both of these concepts to be frequently examined in papers employing LIWC in software engineering. However, contrary to the observations by Lenberg et al., we found that Organizational Climate and Positive Psychology also to be frequently examined and studied in a number of papers. Figure 3 illustrates the most commonly studied BSE concepts in our selected literature.
Figure 3: Most Common BSE Concepts Studied in Software Engineering Literature Using LIWC.
While some BSE concepts, such as sentiment or emotion, have long been studied using a range of automated tools, others like Organizational Climate have traditionally relied on qualitative methods, such as interviews or surveys, due to their abstract and contextual nature. However, LIWC has enabled researchers to examine these social concepts in a more scalable and automated fashion through text analysis. This may help explain why Organizational Climate appears more frequently in our set of studies than reported in earlier BSE reviews [89]. Similarly, BSE and social concepts like Stress and trust, which might otherwise be difficult to quantify or observe without self-reporting, become more accessible through LIWC’s built-in stress-related categories. In this way, LIWC lowers the barrier to analyzing many complex psychological or social phenomena in SE, helping to expand the reach of BSE research.
@@
[backgroundcolor=lightgray!20,topline=false,leftline=false,rightline=false,bottomline=false] RQ3 Summary: We present all the BSE concepts studied with the use of LIWC in Table LABEL:tab:bse. The most commonly examined concepts include Communication, Motivation, Organizational Climate, Positive Psychology, and Personality, while concepts such as Risk Taking, Stress, and Group Dynamics, are less commonly studied.
@@ 5.4 RQ4: How often and in what way has LIWC been evaluated in a software engineering context?
Like any automated analysis tool, LIWC must be evaluated in order for us to determine its effectiveness within the context of SE studies. In the context of our review, evaluation refers to any effort by SE researchers to assess the effectiveness or reliability of LIWC in their studies. Such evaluations could involve expert assessments, for example, a psychologist reviewing whether LIWC is accurately capturing relevant psycholinguistic signals, or comparisons between LIWC output and established ground truth data. To answer RQ4, we identify and categorize the various ways in which software engineering researchers have evaluated the efficacy of LIWC in their studies.
A notable observation among the papers that used LIWC is that the majority, specifically 26 of 43 papers, did not formally evaluate LIWC’s output (see Table 8 and figure 4). These papers relied on LIWC for various purposes and accepted the resulting psychometric measures without any evaluation. In these studies, LIWC’s output was either used for qualitatively analyzing the text or was combined with other data to facilitate subsequent analysis. Some of these studies have used the combination of LIWC’s output and other data points to train machine learning models for tasks such as the prediction of the low quality Stack Overflow posts [S1, S39] or the identification of different arguments within a discussion i.e, argument extraction [S11]. These papers mainly justified their decision based on the established reputation of LIWC as a widely recognized and reliable psychometric text analysis tool, cited by thousands of studies  [22].
@@
Figure 4: Distribution of evaluation strategies used in SE studies employing LIWC.
@@
One study, [S9], performed a comparison of LIWC’s sentiment analysis performance with another model, NLTK VADER, and found them to be perform comparably. However, this was not a formal evaluation of LIWC itself, as it simply compared the performance between the two tools without directly assessing LIWC’s internal accuracy or outputs in isolation. Despite this, we still wanted to distinguish it as a valuable comparison for understanding how LIWC performs relative to other sentiment analysis tools.